In a certain economy marginal propensity to save is 0.2 and the autonomous consumption is 400
a)formulate consumption function
b)if the government expenditure is to increase were to increase by 50% what would be the resultant change in national income
1
Expert's answer
2018-03-28T11:49:08-0400
In a certain economy marginal propensity to save is s = 0.2, c = 1 - s = 0.8 and the autonomous consumption is Ca = 400
a) consumption function is C = Ca + c(Y - T) = 400 + 0.8(Y - T)
b) multiplier m = 1/(1 - c) = 1/(1 - 0.8) = 5.
If the government expenditure were to increase by 50% then the resultant change in national income is 50%*5 = 250%.